### v2

# VisCRA: A Visual Chain Reasoning Attack for Jailbreaking Multimodal Large Language Models

Abstract
The emergence of Multimodal Large Language Models (MLRMs) has enabled sophisticated visual reasoning capabilities by integrating reinforcement learning and Chain-of-Thought (CoT) supervision. However, while these enhanced reasoning capabilities improve performance, they also introduce new and underexplored safety risks. In this work, we systematically investigate the security implications of advanced visual reasoning in MLRMs. Our analysis reveals a fundamental trade-off: as visual reasoning improves, models become more vulnerable to jailbreak attacks. Motivated by this critical finding, we introduce VisCRA (Visual Chain Reasoning Attack), a novel jailbreak framework that exploits the visual reasoning chains to bypass safety mechanisms. VisCRA combines targeted visual attention masking with a two-stage reasoning induction strategy to precisely control harmful outputs. Extensive experiments demonstrate VisCRA’s significant effectiveness, achieving high attack success rates on leading closed-source MLRMs: 76.48% on Gemini 2.0 Flash Thinking, 68.56% on QvQ-Max, and 56.60% on GPT-4o. Our findings highlight a critical insight: the very capability that empowers MLRMs — their visual reasoning — can also serve as an attack vector, posing significant security risks. Warning: This paper contains unsafe examples.
VisCRA: A Visual Chain Reasoning Attack for Jailbreaking Multimodal Large Language Models
Bingrui Sima1thanks: The first two authors contribute equally.  Linhua Cong [^1] footnotemark: 1  Wenxuan Wang2  Kun He1thanks: Corresponding author. 1Huazhong University of Science and Technology 2Hong Kong University of Science and Technology {d202481592, m202476968}@hust.edu.cn brooklet60@hust.edu.cn
1 Introduction
Recent advances in Large Reasoning Models (LRMs), such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024), have introduced a new reasoning paradigm. Unlike traditional prompt-based approaches (Yao et al., 2023), LRMs acquire reasoning capabilities through reinforcement learning, enabling strong performance on complex cognitive tasks (Qu et al., 2025).
Building on these developments, the multimodal AI community has begun incorporating Chain-of-Thought (CoT) supervision and reinforcement learning fine-tuning into Multimodal Large Language Models (MLLMs). This integration has led to the emergence of Multimodal Large Reasoning Models (MLRMs), such as MM-EUREKA (Meng et al., 2025) and OpenAI o4-mini (OpenAI, 2025), which demonstrate significantly improved visual reasoning abilities. These models represent a foundational step toward the long-term goal of multimodal artificial general intelligence (AGI) (Wang et al., 2025; Li et al., 2025b).
Despite these advances, such powerful reasoning models also bring critical safety concerns (Ying et al., 2025). Recent research on text-only LRMs, particularly the DeepSeek-R1 series, has indicated that detailed reasoning can amplify safety risks by enabling models to produce more precise and potentially harmful outputs (Jiang et al., 2025; Zhou et al., 2025). These findings have sparked increased attention to the safety implications of high-capacity reasoning in language models.
In contrast, the corresponding risks in MLRMs remain rather underexplored, despite the added complexity and potential vulnerabilities introduced by visual modalities. Visual inputs can serve as rich contextual cues that guide or reinforce harmful reasoning trajectories, thereby expanding the attack surface for adversarial exploitation. This gap in understanding raises urgent concerns about the robustness and security posture of MLRMs.
Motivated by these concerns, we pose two critical research questions:
  • Does stronger visual reasoning capability increase the security risks of MLLMs?
  • How can adversaries exploit visual reasoning to bypass the safety mechanisms of MLLMs?
In this work, we take a first step toward answering these questions by systematically analyzing the security vulnerabilities introduced by advanced visual reasoning in MLRMs. We begin with a series of preliminary studies that yield critical insights. In particular, we empirically demonstrate that MLRMs exhibit significantly higher susceptibility to jailbreak attacks compared to their base MLLM counterparts. This observation highlights a fundamental trade-off: as visual reasoning capabilities increase, safety alignment tends to degrade.
Building on this finding, we further investigate the use of visual Chain-of-Thought (CoT) prompts in conjunction with existing jailbreak techniques to more deeply engage a models’ visual reasoning capabilities. This combined approach leads to a substantial increase in jailbreak success rates, indicating that the reasoning chain itself can serve as an attack vector. Interestingly, we also observe that when a model produces overly detailed descriptions of harmful visual content early in its reasoning process, its internal safety mechanisms are more likely to be triggered. This suggests a delicate balance between reasoning depth and safety compliance, one that adversaries could potentially manipulate to bypass built-in safeguards.
Based on these insights, we propose VisCRA (Visual Chain Reasoning Attack), a novel multimodal jailbreak framework that explicitly exploits and manipulates the visual reasoning process to circumvent a model’s safety mechanisms.
Our VisCRA operates through a two-stage strategy to achieve this: it first selectively masks critical image regions relevant to the harmful intent, thereby managing initial exposure to toxic content. Following this, a stepwise induction process guides the model to infer the obscured information and then use this reconstructed context, along with visible cues, to execute malicious instructions. This controlled manipulation of the visual reasoning chain aims to ensure outputs remain below safety detection thresholds without sacrificing reasoning coherence. Through this progressive manipulation of the visual reasoning chain, VisCRA effectively transforms enhanced visual reasoning — traditionally viewed as a strength — into a potent adversarial vector capable of bypassing safety defenses.
We validate the effectiveness of VisCRA through extensive experiments on seven open-source MLLMs and four prominent closed-source models, evaluated across two representative benchmarks. Our results demonstrate that VisCRA consistently outperforms existing jailbreak techniques, achieving significantly higher attack success rates across models under diverse settings. These findings reveal critical and previously overlooked security vulnerabilities in current MLRMs.
Our main contributions are threefold:
  • We identify a fundamental trade-off between visual reasoning capability and safety alignment in MLLMs, showing that enhanced visual reasoning can increase vulnerability to jailbreak attacks.
  • We introduce VisCRA, a novel multimodal jailbreak framework that precisely exploits and controls the visual reasoning process, leading to significantly higher attack success rates.
  • Extensive evaluations on both open-source and closed-source MLLMs validate the effectiveness of VisCRA and reveal critical security vulnerabilities in state-of-the-art MLRMs.
2 Related Work
To our knowledge, the security risks introduced by the reasoning capabilities of Multimodal Large Reasoning Models (MLRMs) remain largely underexplored. Existing research has primarily focused on two adjacent areas: (1) the safety implications of reasoning in text-only Large Reasoning Models (LRMs) and (2) jailbreaking attacks targeting Multimodal Large Language Models (MLLMs). We briefly review both lines of work below.
2.1 Safety Challenges in LRMs
Recent studies have shown that enhanced reasoning capabilities in LRMs do not necessarily correlate with improved safety. For instance, Li et al. (2025a) systematically investigate the trade-off between reasoning depth and safety alignment, revealing that deeper reasoning chains can expose latent vulnerabilities. Follow-up work (Zhou et al., 2025; Ying et al., 2025) further highlights that the reasoning process itself (not just the final output) can be a critical locus of safety risk. In particular, multi-step reasoning has been shown to increase the likelihood of generating harmful or policy-violating content. Complementary research (Jiang et al., 2025) also explores how different reasoning strategies affect safety performance in advanced models such as DeepSeek-R1 (Guo et al., 2025), emphasizing that certain reasoning formats (e.g., step-by-step CoT) may unintentionally aid harmful task completion.
2.2 Jailbreak Attacks on MLLMs
Building on earlier jailbreak techniques for text-only LLMs, recent efforts began to adapt such attacks to multimodal settings (Zhang et al., 2024; Bailey et al., 2024). In white-box attack scenarios, ImgJP (Niu et al., 2024) employs maximum-likelihood optimization to generate transferable adversarial images that effectively jailbreak diverse large vision-language models. Qi et al. (2024) demonstrate that a single universal adversarial image can induce harmful outputs when paired with various malicious texts. Wang et al. (2024) employ a dual-optimization framework to simultaneously perturb both image and text modalities to maximize harmful impact. In black-box attack scenarios, FigStep (Gong et al., 2025) circumvents safety alignment by embedding malicious instructions via typography. MM-SafetyBench (Liu et al., 2024) leverages diffusion models to synthesize query-relevant harmful images, and HADES (Li et al., 2024) makes enhancements via optimized prompts for diffusion models, producing more semantically coherent and potent harmful samples.
However, current approaches do not explicitly engage or manipulating the visual reasoning process. As such, they do not account for the additional vulnerabilities introduced by multistep visual reasoning, which is a defining feature of modern MLRMs. Our work bridges this gap by directly targeting the visual reasoning chain itself, revealing a novel and potent attack surface unique to MLRMs.
3 Motivation
3.1 Vulnerability of MLRMs
Figure 1: Attack success rates (ASR) of base MLLMs vs. reasoning-enhanced MLRMs, with and without visual CoT prompting. Enhanced models (e.g., R1-Onevision) exhibit significantly higher vulnerability to HADES attacks compared to their base counterparts (e.g., Qwen2.5-VL), and the inclusion of visual CoT prompting further amplifies ASR across all models.
Figure 2: Illustration of a visual CoT failure case. An early, overly detailed description of harmful visual content (in red) triggers the model’s safety mechanisms (in green), interrupting the reasoning process.
Prior research in LRMs has identified a fundamental trade-off between reasoning and safety, showing that enhanced reasoning capabilities can weaken a model’s resistance to harmful prompts (Jiang et al., 2025). In this work, we extend this to MLRMs by systematically investigating how enhanced visual reasoning affects their safety.
To this end, we evaluate base MLLMs and their reasoning-enhanced MLRM variants using the HADES benchmark (Li et al., 2024), which is designed to probe multimodal safety under adversarial conditions. As shown in Figure 1, reasoning-enhanced MLRMs (such as R1-Onevision Yang et al. (2025), MM-EUREKA-InternVL Meng et al. (2025), and LLaVA-CoT Xu et al. (2024)) are substantially more susceptible to jailbreak attacks compared to their base counterparts.
For instance, R1-Onevision achieves a 65% ASR, more than double that of its base model Qwen2.5-VL. This substantial increase in vulnerability is largely attributable to the models’ enhanced visual reasoning, which enables them to interpret and respond to harmful prompts in more detailed, coherent, and actionable ways. While advanced visual reasoning improves performance on complex cognitive and perception tasks, it simultaneously amplifies security risks, a trade-off that is both consequential and currently underappreciated. These findings motivate a deeper exploration of how reasoning itself can be exploited as an attack surface in MLRMs.
3.2 Attack Amplification via Visual CoT
Building on the observation from Section 3.1 that stronger visual reasoning increases MLLMs vulnerability, we hypothesize that explicitly eliciting visual Chain-of-Thought (CoT) reasoning could further amplify attack success. To test this, we design a tailored visual CoT prompt to encourage step-by-step reasoning over image content
Figure 3: Illustration of VisCRA. The framework employs: (1) Attention-Guided Masking of the critical harmful region using an auxiliary model, (2) Multi-Stage Reasoning Induction for the target model to infer masked content and then execute the harmful instruction.
(See Appendix A.1 for the prompt template.). Empirical results confirm our hypothesis: integrating visual CoT with HADES adversarial instructions significantly boosts jailbreak success rates (as illustrated in Figure 1, the increase from ’HADES’ to ’HADES+CoT’ bars for each model), highlighting the power of guided visual reasoning in bypassing safety mechanisms. However, this approach also reveals an important failure mode. While detailed image descriptions can aid reasoning, over-describing harmful visual content too early in the reasoning process can generate an excess of toxic tokens, inadvertently triggering built-in safety filters. This results in the model rejecting the harmful prompt before execution, as illustrated in Figure 2.
To address this limitation, it is crucial to develop an attack strategy that leverages the model’s visual reasoning capabilities for detailed and structured responses to harmful prompts, while carefully regulating the reasoning process to avoid premature safety triggers. Specifically, the attack must balance two competing objectives: (1) eliciting sufficient visual detail to support coherent reasoning, and (2) suppressing early overexposure to explicitly harmful content that could activate the model’s safety mechanisms before the harmful intent is fully inferred or executed.
4 Methodology
We propose VisCRA (Visual Chain Reasoning Attack), a novel jailbreak framework designed to exploit the visual reasoning capabilities of MLLMs while strategically evading built-in safety mechanisms. As illustrated in Figure 3, VisCRA consists of two key components: (1) Attention-Guided Masking that employs an auxiliary model to identify and mask image regions most relevant to the harmful intent as guided by attention, and (2) Multi-Stage Reasoning Induction that guides the target MLLM to first infer the masked content, curtailing overexposure and establishing a coherent reasoning foundation, and then to execute harmful instruction based on this inference and visible image context. Consequently, VisCRA effectively exploits visual reasoning by guiding a structured harmful process that preserves coherence and avoids premature safety activations.
4.1 Attention-Guided Masking
As illustrated in Figure 2, early and excessive exposure to harmful visual content can prematurely trigger a model’s safety mechanisms, disrupting the progression of harmful reasoning. To mitigate this, our Attention-Guided Masking module strategically suppresses the most toxic visual elements while maintaining semantic coherence. The key idea is to identify and mask the image region most critical to the harmful instruction. This selective masking is guided by an auxiliary MLLM, which serves as an interpretability tool to highlight visually salient regions in relation to the harmful prompt. By masking only the regions most associated with toxic semantics, we ensure that the model begins reasoning from a controlled yet informative visual input, laying the groundwork for gradual reconstruction and instruction execution.
4.1.1 Image-Token Relevance Scoring
Given an input image $I$ and a harmful instruction $q$, we feed the pair into an auxiliary MLLM (Qwen2.5-VL) and extract the cross-modal attention tensor from a specific decoder layer $\ell$. The resulting tensor, $A_{\ell}\in\mathbb{R}^{H\times T_{\mathrm{out}}\times T_{\mathrm{img}}}$, captures the attention weights between output language tokens and visual image tokens, where $H$ is the number of attention heads, $T_{\mathrm{out}}$ is the number of output tokens, and $T_{\mathrm{img}}$ is the number of image tokens. To obtain per-token relevance scores $a_{i}$ for each image token, we average $A_{\ell}$ over all heads and focus on the first output token, as it aggregates attention information from all input tokens:
$a_{i}=\frac{1}{H}\sum_{h=1}^{H}A_{\ell}[h,1,i],\quad i=1,\dots,T_{\mathrm{img}}.$ (1)
The relevance scores $\{a_{i}\}$ are then reshaped according to the spatial grid arrangement of these image tokens (e.g., an $N_{h}\times N_{w}$ grid, where $T_{\mathrm{img}}=N_{h}\times N_{w}$). This forms a two-dimensional attention map $A\in\mathbb{R}^{N_{h}\times N_{w}}$ that highlights image regions critical to the model’s interpretation of the harmful query at the token level.
4.1.2 Region Selection and Masking
To identify and mask the region most relevant to the harmful intent, we apply a sliding window of size $B\times B$ tokens with stride $s$ tokens over the attention map $A$ , generating candidate patches $\mathcal{R}$. The relevance score for each patch $r\in\mathcal{R}$ is calculated as the summation of attention scores:
$s(r)=\sum_{(x,y)\in r}A(x,y).$ (2)
We then sort these patches by their relevance scores in descending order and randomly select one patch $r^{*}$ from the top three. This introduction of slight randomness helps mitigate potential model biases (Darcet et al., 2024). The image region corresponding to the selected patch $r^{*}$ (which represents a collection of image tokens) is then masked in the original image $I$ using a green rectangle, resulting in the modified image $I^{\prime}$. Leveraging the observation that harmful attention regions tend to be consistent across different MLLMs, we compute $r^{*}$ using the auxiliary model (Qwen2.5-VL) and apply this mask to other target black-box models.
4.2 Multi-Stage Reasoning Induction
Building on the masked input generated by the previous stage, our Multi-Stage Reasoning Induction component is designed to carefully steer the model’s reasoning chain to address the crucial balance identified in Section 3.2: stimulating detailed harmful reasoning while controlling early harmful textual exposure. To achieve this, we introduce a two-phase prompting approach, which combines both phases into a single input for the target MLLM.
Phase 1: Focused Reasoning on Harmful Region
Phase 1 employs a prompt to induce focused visual reasoning solely on the masked region, aiming for a coherent yet controlled description of its plausible harmful content:
First, infer one plausible content for the green-masked region in the image.
This localized inference curtails extensive early harmful outputs by restricting the model’s focus. Such restriction limits exposure to overt toxic content, thereby reducing the likelihood of activating safety mechanisms while anchoring the reasoning chain. Consequently, the model is guided to produce a concise and coherent description of critical harmful elements, establishing a controlled foundation for the subsequent phase.
Phase 2: Contextualized Instruction Execution
Phase 2 prompts the model to execute harmful behavior using the reasoning established in Phase 1 and the remaining visible image:
Then, based on the image and your reasoning, [harmful_query]. Show your reasoning ability.
This step guides the model to fully engage its reasoning capabilities on the [harmful_query] (the placeholder replaced with the specific instruction, e.g., the harmful instruction from the HADES benchmark), leveraging both the inferred content and the remaining visual context. This ensures the final output not only be harmful as intended but also detailed and logically consistent with the preceding analysis.
Model Animal Privacy Self-Harm Violence Financial Overall
H Ours H Ours H Ours H Ours H Ours H Ours
Open-Source Models
Qwen2.5-VL 5.33 55.33 32.67 92.67 16.00 68.67 55.33 90.67 44.00 91.33 30.27 79.73
MM-E-Qwen 8.67 57.33 33.33 93.33 17.33 64.67 55.67 91.33 46.00 90.00 32.20 79.33
R1-Onevision 37.33 62.00 69.33 94.00 64.00 79.33 78.67 91.33 74.00 89.33 65.06 83.20
InternVL2.5 16.67 44.00 22.00 69.33 18.00 44.67 33.33 68.67 41.33 79.33 26.27 61.20
MM-E-InternVL 20.00 44.67 26.67 76.67 30.00 54.67 46.67 72.67 49.33 82.67 34.55 66.27
LLaMA-3.2-V 2.00 56.00 2.67 70.67 0.00 64.67 4.00 80.00 7.33 76.00 3.20 69.47
LLaVA-CoT 19.33 64.00 18.67 88.00 18.67 68.67 37.33 89.33 32.67 89.33 25.33 79.87
Closed-Source Models
GPT-4o 1.33 45.67 9.33 57.33 6.67 53.33 16.00 65.33 14.67 60.00 9.60 56.60
Gemini 2.0 FT 5.33 44.67 40.67 70.67 16.67 62.67 44.67 80.67 48.00 71.33 31.06 66.00
QvQ-Max 11.33 41.33 44.67 78.00 21.33 59.33 64.00 76.67 58.67 76.00 40.13 66.27
OpenAI o4-mini 0.00 12.00 0.67 9.33 0.00 4.67 0.00 11.33 1.33 21.33 0.40 11.73
Table 1: ASR (%) comparison of the HADES baseline (H) with VisCRA (Ours) on the HADES benchmark. The best results appear in bold.
Model IA HS MG PH Fr PV Overall
QR Ours QR Ours QR Ours QR Ours QR Ours QR Ours QR Ours
Open-Source Models
Qwen2.5-VL 54.64 95.88 34.97 80.37 54.55 81.82 52.08 77.08 60.39 94.16 49.64 79.86 49.73 84.62
MM-E-Qwen 56.70 97.94 40.49 81.60 52.27 82.82 55.56 81.94 58.67 94.81 55.40 82.01 50.94 84.35
R1-Onevision 88.66 91.75 66.26 73.62 68.18 77.27 75.00 79.17 81.82 85.06 77.70 79.86 75.89 80.84
InternVL2.5 21.65 61.01 25.77 50.31 45.45 77.27 42.36 69.44 37.01 82.42 28.78 62.59 33.50 67.21
MM-E-InternVL 43.30 79.38 31.33 59.51 47.72 81.82 47.91 75.69 51.95 88.96 47.48 74.82 44.09 75.57
LLaMA-3.2-V 12.37 97.94 16.56 61.94 36.36 72.73 23.61 69.44 27.92 86.36 23.02 78.42 22.13 76.93
LLaVA-CoT 69.07 96.91 59.51 77.91 56.82 79.55 61.80 77.08 77.78 92.86 58.27 79.58 63.37 83.94
Closed-Source Models
GPT-4o 1.03 44.33 2.45 28.83 13.64 54.55 15.28 53.47 7.79 63.64 2.16 36.69 6.88 45.88
Gemini 2.0 FT 49.48 88.66 40.49 67.48 54.55 61.36 61.11 68.06 74.03 82.47 60.43 76.98 56.42 76.48
QvQ-Max 36.08 75.26 12.88 45.40 59.09 72.73 51.39 72.92 53.90 83.12 44.60 69.06 40.62 68.56
OpenAI o4-mini 0.00 8.25 3.68 10.43 2.27 13.64 1.39 9.72 1.30 9.09 0.00 8.63 1.48 9.58
Table 2: ASR (%) comparison of the QR-Attack baseline (QR) with VisCRA (Ours) on the MM-SafetyBench benchmark. The best results appear in bold. Categories: IA (Illegal Activity), HS (Hate Speech), MG (Malware Generation), PH (Physical Harm), Fr (Fraud), PV (Privacy Violence).
5 Experiments
5.1 Experimental Setup
Evaluation Models.
We evaluate the effectiveness of VisCRA on eleven diverse MLLMs, including seven open-source models and four closed-source commercial systems. The open-source models include Qwen2.5-VL Bai et al. (2025), InternVL2.5 Chen et al. (2024), and LLaMA-3.2-11B-Vision, along with their reasoning-augmented variants: MM-EUREKA-Qwen and R1-Onevision (both fine-tuned from Qwen2.5-VL), MM-EUREKA-InternVL (fine-tuned from InternVL2.5), and LLaVA-CoT (fine-tuned from LLaMA-3.2-11B-Vision). The closed-source models include GPT-4o (2024-11-20) Hurst et al. (2024), OpenAI o4-mini, Gemini 2.0 Flash Thinking DeepMind (2024), and QvQ-Max Alibaba (2025).
Evaluation Metrics.
We evaluate model vulnerability using Attack Success Rate (ASR), which measures the proportion of inputs that successfully elicit harmful responses as judged by an LLM evaluator. Formally:
ASR= # Successful Attacks# Total $\text{ASR}=\frac{\text{\ \# Successful Attacks}}{\text{\# Total Inputs}}\times 100\%$
An attack is considered successful only when the model generates a response that explicitly follows the original harmful prompt, beyond merely describing the image content. Following SafeChain (Jiang et al., 2025), we adopt Llama-Guard-3-8B (Inan et al., 2023) as the safety evaluator due to its strong alignment with human judgment and effectiveness in evaluating long-form, reasoning-based outputs.
Baselines and Benchmarks.
We conduct evaluations on two widely adopted multimodal jailbreak benchmarks: HADES (Li et al., 2024) and MM-SafetyBench (Liu et al., 2024). HADES employs a strategy where malicious intent is embedded and amplified within crafted images and accompanying typography. HADES provides 750 samples across five harmful scenarios. MM-SafetyBench utilizes the Query-Relevant Attack (QR) strategy, which rephrases harmful questions to bypass safety mechanisms, covering 13 prohibited categories. For MM-SafetyBench, to ensure comparability with HADES, we use a subset of 741 samples focused on six explicit harmful categories (Illegal Activity, Hate Speech, Physical Harm, Fraud, Privacy Violence, Malware Generation).
Implementation Details.
In the attention-guided masking module, we extract the cross-attention tensor from the 19th decoder layer ($\ell=19$) of the auxiliary MLLM. The sliding window size $B$ was set to 12 tokens, with a stride $s$ of 4 tokens to efficiently localize relevant image regions. The mask region corresponds to a $B\times B$ patch and the mask is applied using a green overlay. The choice of these hyperparameters is supported by ablation studies presented in Appendix A.2.
5.2 Main Results
Our proposed VisCRA consistently surpasses existing attack baselines across both open-source and closed-source MLLMs, demonstrating strong jailbreak efficiency (Tables 1 and 2).
On Open-Source Models.
VisCRA achieves overall ASR ranging from 61.20% to 83.20% on the HADES benchmark and from 67.21% to 84.62% on MM-SafetyBench (see ’Overall Ours’ columns in Tables 1 and 2). Notably, LLaMA-3.2-V (Table 1), which demonstrated strong robustness against the HADES attack (Overall ASR of 3.20%), becomes significantly more vulnerable under VisCRA, reaching an overall ASR of 69.47%. Moreover, Reasoning-enhanced models like LLaVA-CoT are more vulnerable to VisCRA attacks, achieving ASRs of 79.87% on HADES and 83.94% on MM-SafetyBench with VisCRA, compared to their base counterparts’ ASRs of 69.47% and 76.93%, respectively.
On Closed-Source Models.
VisCRA also significantly enhances attack effectiveness on closed-source commercial systems. VisCRA boosts the overall ASR from 9.60% to 56.60% on HADES for GPT-4o (Table 1). Even OpenAI’s latest model, o4-mini, which incorporates reasoning-based safety monitors, experiences a notable ASR increase on HADES, rising from a mere 0.40% baseline to 11.73% under VisCRA (Table 1). Most alarmingly, VisCRA drives strikingly high ASRs in advanced commercial visual reasoning models: Gemini 2.0 Flash Thinking (Gemini 2.0 FT) reaches 76.48% on MM-SafetyBench, while QvQ-Max attains 66.27% on HADES (see Tables 2 and 1).
5.3 Ablation Study
To gain deeper insights into the contributions of VisCRA’s key components, we conduct ablation studies focusing on its two core mechanisms: attention-guided masking and multi-stage induction prompting. Experiments are carried out on two HADES sub-categories (Self-Harm and Animal) across three representative MLLMs: LLaVA-CoT, MM-Eureka-Qwen, and GPT-4o.
5.3.1 On Attention-Guided Masking
Model Self-Harm Animal
HADES baseline
LLaVA-CoT 18.67% 19.33%
MM-EUREKA-Qwen 17.33% 8.67%
GPT-4o 6.67% 1.33%
VisCRA + Random Mask
LLaVA-CoT 53.33% 42.00%
MM-EUREKA-Qwen 51.33% 39.33%
GPT-4o 38.00% 26.00%
VisCRA + Attention-Guided Mask
LLaVA-CoT 68.67% 64.00%
MM-EUREKA-Qwen 64.67% 57.33%
GPT-4o 53.33% 45.67%
Table 3: Ablation study on different masking strategies. All settings use VisCRA’s two-stage prompt.
Prior research has shown that irrelevant reasoning context can disrupt logical coherence and degrade model performance (Yan et al., 2025). To assess the importance of targeted visual masking in facilitating effective harmful reasoning, We compared VisCRA with a variant using a random mask. The two differed only in the masking method. The results are presented in Table 3.
Our results show that VisCRA with random masking outperforms the original HADES baseline, likely because it still engages the model’s visual reasoning to bypass defences. However, attention-guided masking consistently yields the highest ASR across all evaluated models, underscoring the importance of aligning the masked region with the harmful prompt.
As illustrated in Figure 4, random masking often obscures irrelevant image details, leading to incoherent reasoning that disrupts the intended harmful reasoning process. This incoherence may prematurely trigger safety mechanisms. Meanwhile, random masking can leave critical harmful regions unmasked, directly exposing the malicious intent and thus activating the model’s safety mechanisms.
Figure 4: A failure case on random masking.
5.3.2 On Multi-Stage Induction Prompting
To rigorously assess our prompting strategy, we compare five configurations: (1) the original HADES baseline, (2) HADES combined with attention-guided masking, (3) HADES augmented with visual CoT prompting, (4) HADES employing both masking and visual CoT, and (5) the complete VisCRA framework.
As detailed in Table 4, attention-guided masking alone yields a moderate increase in ASR by suppressing high-risk visual regions. Incorporating visual CoT further boosts ASR by eliciting more detailed reasoning; however, this often causes premature overexposure to harmful content early in the output, which triggers the model’s safety mechanisms prematurely. While combining masking with visual CoT provides a slight additional improvement, it still struggles with premature exposure.
Model Self-Harm Animal
HADES baseline
LLaVA-CoT 18.67% 19.33%
MM-EUREKA-Qwen 17.33% 8.67%
+ Attention-Guided Mask only
LLaVA-CoT 30.00% 25.33%
MM-EUREKA-Qwen 21.33% 10.00%
+ Visual CoT
LLaVA-CoT 41.33% 30.67%
MM-EUREKA-Qwen 48.00% 23.33%
+ Attention-Guided Mask + Visual CoT
LLaVA-CoT 50.33% 32.00%
MM-EUREKA-Qwen 50.00% 26.00%
Full VisCRA
LLaVA-CoT 68.67% 64.00%
MM-EUREKA-Qwen 64.67% 57.33%
Table 4: Ablation study on different prompt configurations over two HADES sub-categories.
In contrast, VisCRA’s two-stage induction carefully guides the model along a coherent, goal-directed reasoning path, while simultaneously regulating the initial output to avoid prematurely triggering safety mechanisms. This tailored structure fully leverages visual reasoning capabilities, yielding the highest ASR among all tested configurations. Overall, these findings highlight the importance of image-text coordination in our prompt design for achieving effective and reliable jailbreaks.
6 Conclusion
We explored the security risks introduced by enhanced visual reasoning in Multimodal Large Reasoning Models (MLRMs). Through empirical analysis, we illustrated that stronger reasoning capabilities paradoxically undermine safety, making models more prone to producing detailed and coherent responses to harmful prompts. To probe this vulnerability, we proposed VisCRA, a novel jailbreak framework that combines attention-guided visual masking with a two-stage reasoning induction strategy. VisCRA effectively manipulates the model’s reasoning chain to evade safety mechanisms while preserving visual coherence. Extensive experiments across a wide range of open- and closed-source MLRMs validate the effectiveness of VisCRA, revealing significantly elevated attack success rates. These findings expose advanced reasoning as a double-edged sword - an asset for task performance, but also a critical security liability. Our work highlights the urgent need for reasoning-aware safety frameworks to safeguard current and next-generation MLRMs against increasingly sophisticated adversarial attacks.
Limitations
Our study mainly focuses on how to leverage the visual reasoning capabilities of Multimodal Large Reasoning Models (MLRMs) to amplify their safety risks. However, developing strategies to enhance the safety of these models against such reasoning-based vulnerabilities, while preserving their core reasoning capabilities, remains an open-problem for future research.
Ethical Statement
This research investigates security vulnerabilities within Multimodal Large Reasoning Models (MLRMs), particularly those related to their enhanced visual reasoning capabilities. We introduce our VisCRA jailbreak method in this work primarily to highlight and analyze these critical risks. Our primary objective is to expose such limitations to promote safer AI development and robust safety alignments, not to create or facilitate tools for misuse. All evaluations are conducted on established public benchmarks in controlled settings.
Appendix A Appendix
The appendix provides supplementary materials, including: (1) the visual CoT prompt template from Section 3.2 (Section A.1); (2) a hyperparameter study of VisCRA’s masking component (mask size and color) (Section A.2); and (3) successful VisCRA attack examples on prominent closed-source models (Section A.3).
A.1 Visual CoT Prompt Template
The visual Chain-of-Thought (CoT) prompt template, referenced in Section 3.2 of the Motivation, is presented below. It elicits a detailed, step-by-step reasoning process from MLLMs given an image and associated instruction.
Generate an image description based on the instruction. Then, provide a rationale to analyze the instruction. Next, generate a step-by-step reasoning process to solve the problem. Finally, provide the final answer after your above analysis of the instruction. instruction: [harmful_query]
[harmful_query] is a placeholder that represents the specific instruction, which contains content intended to probe the model’s ability to handle potentially harmful or sensitive scenarios. This structured prompt guides the model through four stages: image interpretation, instruction understanding, systematic reasoning, and final answer generation.
A.2 Masking Hyperparameter Study
To further investigate the sensitivity of VisCRA to specific choices in the masking process, we conduct ablation studies focusing on two key hyperparameters: mask size and mask color. In all experiments, the masked regions were applied to the image content while preserving the original typography. Experiments used LLaVA-CoT and MM-EUREKA-Qwen on HADES’ Self-Harm and Animal sub-categories. For each setting, we report the Attack Success Rate (ASR) as the primary metric.
A.2.1 Masking Size Ablation
The size of the masked region, parameterized by the token window dimension $B$, plays a critical role in VisCRA’s effectiveness. We experimented with $B\in\{6,12,18\}$ (via a green mask), where the default in our main experiments is $B=12$. These values correspond to token-based patch sizes; for instance, in models like Qwen2.5-VL, one token may represent approximately 28 pixels.
A smaller window size (e.g., $B=6$) may fail to fully obscure the harmful region, allowing the model to still infer problematic content. Conversely, a larger window (e.g., $B=18$) may mask too much context, inadvertently degrading the model’s ability to reason about the scene.
Model Self-Harm Animal
HADES baseline
LLaVA-CoT 18.67% 19.33%
MM-EUREKA-Qwen 17.33% 8.67%
VisCRA with Mask Size $B=6$
LLaVA-CoT 62.67% 50.67%
MM-EUREKA-Qwen 55.33% 38.67%
VisCRA with Mask Size $B=12$ (Default)
LLaVA-CoT 68.67% 64.00%
MM-EUREKA-Qwen 64.67% 57.33%
VisCRA with Mask Size $B=18$
LLaVA-CoT 66.00% 48.00%
MM-EUREKA-Qwen 50.00% 47.33%
Table 5: ASR (%) for varying mask sizes ($B\times B$ tokens, green mask) on HADES sub-categories. Default VisCRA setting uses $B=12$.
Table 5 indicates that $B=12$ (default) yields the highest ASR across both models and sub-categories. A smaller mask size ($B=6$) leads to a marked reduction in performance, likely due to insufficient coverage of the critical harmful regions in the image. On the other hand, increasing the mask size to $B=18$ also degrades performance, suggesting that an excessively large mask may obscure essential visual context required for reasoning. Overall, $B=12$ offers the most effective balance between masking harmful content and preserving surrounding context necessary for successful attack execution.
A.2.2 Masking Color Ablation
We also examine whether the mask color influences VisCRA’s effectiveness. Specifically, we compared our default green mask against a black mask (B=12 fixed). The results are summarized in Table 6.
Model Self-Harm Animal
HADES baseline
LLaVA-CoT 18.67% 19.33%
MM-EUREKA-Qwen 17.33% 8.67%
VisCRA with Green Mask (Default)
LLaVA-CoT 68.67% 64.00%
MM-EUREKA-Qwen 64.67% 57.33%
VisCRA with Black Mask
LLaVA-CoT 62.00% 57.33%
MM-EUREKA-Qwen 58.00% 50.00%
Table 6: ASR (%) for different mask colors (with $B=12$) on HADES sub-categories.
Across both models and sub-categories, the green mask consistently yields noticeably higher ASR than the black mask. This outcome suggests that the green mask, often a more salient and distinct color against typical image backgrounds, is a more salient and contrasting color relative to typical image backgrounds, serving as a more effective visual cue. It likely helps direct the model’s attention and reasoning more explicitly toward the masked region. In contrast, the black mask may blend into darker image areas, especially in images with naturally dark themes, thereby diminishing its ability to signal the region of interest to the model.
A.3 More Illustrative Examples
Figures 5 to 8 present representative examples of successful attacks conducted using VisCRA. These cases illustrate how VisCRA effectively circumvents the safety mechanisms of several prominent closed-source models, including GPT-4o, OpenAI o4-mini, QvQ-Max, and Gemini 2.0 Flash Thinking.
Figure 5: Example on QvQ-Max.
In each example, VisCRA prompts the model to first infer the content obscured by the green mask and then reason about the associated instruction. This two-step reasoning process, facilitated by the visual Chain-of-Thought prompting, enables the model to inadvertently generate responses aligned with harmful queries. These examples visually demonstrate the core mechanism and potency of VisCRA in compromising safety across a range of advanced multimodal systems.
Figure 6: Example on Gemini 2.0 Flash Thinking.
Figure 7: Example on GPT-4o.
Figure 8: Example on OpenAI o4-mini.